# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 3. Tools
Note that if ${\mathbf{A}},\theta,p$ are as in the statement of Theorem 3.8, then every $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e)$ can be viewed as a vector space over $\mathbb{Z}_{p}$. In fact, Freese’s proof of Theorem 3.8 (in the congruence modular case) actually showed that each $\operatorname{Grp}_{{\mathbf{A}}}(\theta,e_{i})$ carries the structure of a vector space over $\mathbb{F}$. Freese proved something similar when ${\mathbf{A}}$ is infinite: there is a division ring $\mathbb{F}$ determined by ${\mathbf{A}}$ and the abelian minimal congruence $\theta$ such that every $\theta$-class inherits the structure of a left $\mathbb{F}$-vector space. We will show in Section 6 that Freese’s proof carries through assuming only that ${\mathbf{A}}$ has a weak difference term.
@@
Suppose ${\mathbf{A}}$ is an with a weak difference term, $\theta\in\operatorname{Con}({\mathbf{A}})$, and $\theta$ is abelian. Let $D_{1},D_{2}$ be subuniverses of ${\mathbf{A}}$ which are transversals for $\theta$. Then there exists an automorphism $\sigma$ of ${\mathbf{A}}$ satisfying $\sigma(D_{1})=D_{2}$ and $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}x$ for all $x\in A$.
@@
Observe first that $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}\tau(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}d(x,x,x)=x$ for all $x\in A$. Fix a $\theta$-class $B$; by construction, there exist $b,c\in B$ such that $\pi_{1}(x)=b$ and $\pi_{2}(x)=c$ for all $x\in B$. Pick an abelian group operation $+$ on $B$ such that $d(x,y,z)=x-y+z$ for all $x,y,z\in B$. Then $\sigma(x)=x-b+c$ and $\tau(x)=x-c+b$ for all $x\in B$, and since this is true for every $\theta$-class, we get $\sigma\circ\tau=\tau\circ\sigma=\mathrm{id}_{A}$, proving $\sigma$ is a bijection. Our aim is to prove that $\sigma$ is the desired automorphism.
Let $f$ be an $n$-ary basic operation. Let $a_{1},\ldots,a_{n}\in A$ and put $a=f(a_{1},\ldots,a_{n})$; we need to prove $\sigma(a)=f(\sigma(a_{1}),\ldots,\sigma(a_{n}))$. Let $b_{i}=\pi_{1}(a_{i})$ and $c_{i}=\pi_{2}(a_{i})$, and note that $a_{i},b_{i},c_{i}$ belong to a common $\theta$-class $B_{i}$ for each $i=1,\ldots,n$. Let $b=f(b_{1},\ldots,b_{n})$ and $c=f(c_{1},\ldots,c_{n})$, and note that $b\in D_{1}$ and $c\in D_{2}$ (since $D_{1},D_{2}\leq{\mathbf{A}}$), so $\pi_{1}(a)=b$ and $\pi_{2}(a)=c$. Thus what must be shown is
@@
This is a special case of [17, Theorem 6.16]. ∎
@@
(1) follows from [17, Theorem 3.27] as corrected in [18]. (2) is a special case of [17, Theorem 6.24]. ∎
Our last result in this section generalizes [14, Lemma 2.6].
@@
$\delta\wedge\tau=\sigma$ implies $C(\delta,\tau;\sigma)$ by e.g. [17, Theorem 2.19(8)], proving $(\sigma:\tau)\geq\delta$. Thus to prove $(\delta:\varepsilon)=(\sigma:\tau)$, it is enough to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\sigma)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. We have $C(\alpha,\tau;\sigma)\iff C(\alpha,\tau;\delta)$ by [17, Theorem 2.19(4)], since $\tau\wedge\sigma=\tau\wedge\delta$. Thus we only need to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\delta)$. The forward implication follows immediately by monotonicity. Conversely, suppose $C(\alpha,\tau;\delta)$. Then $C(\alpha,\delta\circ\tau\circ\delta;\delta)$ by [17, Theorem 2.19(3)], i.e., $C(\alpha,\varepsilon;\delta)$.
(2) If $\varepsilon/\delta$ is abelian, then we can apply item (1). Assume $\varepsilon/\delta$ is nonabelian. Then $\tau/\sigma$ is also nonabelian by Proposition 3.12(1). As in the proof of (1), we only need to prove $C(\alpha,\tau;\delta)\implies C(\alpha,\varepsilon;\delta)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. Assume $C(\alpha,\tau;\delta)$. If $\alpha\geq\varepsilon$, then $C(\tau,\tau;\delta)$ by monotonicity, which contradicts nonabelianness of $\tau/\sigma$. Thus $\alpha\ngeq\varepsilon$. Since $\alpha\geq\delta$ and $\delta\prec\varepsilon$, we get $\alpha\wedge\varepsilon=\delta$ and hence $C(\alpha,\varepsilon;\delta)$ as required. ∎
4. Generalizing an old construction
In this section we revisit an old construction of Hagemann and Herrmann [11] which they used to characterize the centrality relation in congruence modular varieties. Freese [5] adapted the construction to give a clean characterization of his similarity relation in congruence modular varieties. We will show that the construction adapts nicely to varieties with a weak difference term, and we will use it in Section 7 to define similarity in that setting.
In the following definition, we follow the notation of [6].
@@ 4. Generalizing an old construction
    $\mbox{If~{}}\begin{pmatrix}a&c\\ b&d\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi),~{}\mbox{then}~{}\begin{pmatrix}b&d\\ a&c\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi).\$
  3. (3)
    If ${\mathbf{A}}$ has a weak difference term and $\varphi$ is abelian, then $\Delta_{\mathsf{h}}(\theta,\varphi)=M(\theta,\varphi)$.
@@
(1) $M_{\theta,\varphi}$ is a reflexive, symmetric subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$.
(2) This follows from (1) and the fact that the generating set $X(\theta,\varphi)$ for $M(\theta,\varphi)$ is vertically symmetric.
(3) Let $\rho=M_{\theta,\varphi}$; then $\rho$ is a reflexive subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$. If $\varphi$ is abelian and ${\mathbf{A}}$ has a weak difference term $d$, then $d$ interpreted coordinatewise in ${\mathbf{A}}(\theta)$ satisfies the hypothesis of Lemma 3.1 for $\rho$. Thus $\rho$ is transitive by Lemma 3.1, implying $\Delta_{\theta,\varphi}=M_{\theta,\varphi}$ and thus $\Delta_{\mathsf{h}}(\theta,\varphi)=M(\theta,\varphi)$.
(4) We can consider $\Delta_{\mathsf{h}}(\theta,\varphi)$ as a binary relation on the set of its rows. As such, it is a reflexive subuniverse of ${\mathbf{A}}(\varphi)\times{\mathbf{A}}(\varphi)$, and if $\theta$ is abelian, a weak difference term for ${\mathbf{A}}$, interpreted coordinatewise in ${\mathbf{A}}(\varphi)$, will satisfy the hypothesis of Lemma 3.1 for this relation. ∎
@@
The next theorem has its roots in [11, Proposition 1.1].
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is a nonzero abelian congruence of ${\mathbf{A}}$, and $\alpha=(0:\theta)$. Let $\Delta=\Delta_{\theta,\alpha}$, let $\eta_{1},\eta_{2}$ be the kernels of the two projection maps $\operatorname{pr}_{1},\operatorname{pr}_{2}:{\mathbf{A}}(\theta)\rightarrow{\mathbf{A}}$, and let $\varepsilon=\overline{\theta}\wedge\Delta$. Then:
@@
[…]
@@
As $\overline{\theta}$ is abelian, we use Proposition 3.10(2) to get $C(\overline{\theta},\overline{\theta};\varepsilon)$, i.e., $\overline{\theta}/\varepsilon$ is abelian. As $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get that $\overline{\alpha}/\Delta$ is abelian by Proposition 3.12(1). Since also $\overline{\theta}/\eta_{1}$ is abelian and $(\eta_{1},\overline{\theta})\searrow(0,\eta_{2})\nearrow(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get $(\eta_{1}:\overline{\theta})=(0:\eta_{2})=(\varepsilon:\overline{\theta})=(\Delta:\overline{\alpha})$ by Proposition 3.13(1). Finally, $(\eta_{1}:\overline{\theta})=\overline{\alpha}$ by Lemma 2.3(2) and our choice of $\alpha$, so $(\Delta:\overline{\alpha})=\alpha$. This completes the proof of (2).
Now assume $0\prec\theta$. Then $\eta_{i}\prec\overline{\theta}$ for $i=1,2$. Since $\eta_{1}\wedge\eta_{2}=0$, we get $0\prec\eta_{1},\eta_{2}$ (see e.g. [26, Corollary 2.29(ii)]), and every maximal chain in the interval $I[0,\overline{\theta}]$ has length 2 [26, Theorem 2.37]. Thus $I[0,\overline{\theta}]$ has height 2. Note that this also implies $0\prec\varepsilon\prec\overline{\theta}$. Because $\overline{\theta}/\varepsilon$ is abelian and $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, the map $J:I[\varepsilon,\overline{\theta}]\rightarrow I[\Delta,\overline{\alpha}]$ given by $J(\gamma)=\gamma\vee\varepsilon$ is surjective by Proposition 3.12(2). Thus $\Delta\prec\overline{\alpha}$. Now let $\varphi\in\operatorname{Con}({\mathbf{A}}(\mu))$ with $\Delta\leq\varphi$ but $\overline{\alpha}\nleq\varphi$. Then $\overline{\alpha}\wedge\varphi=\Delta$, so $C(\varphi,\overline{\alpha};\Delta)$, so $\varphi\leq(\Delta:\overline{\alpha})=\overline{\alpha}$, so $\varphi=\overline{\alpha}\wedge\varphi=\Delta$. This proves (6). ∎
The next definition is from [5, 6] in the congruence modular setting.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\theta\in\operatorname{Con}({\mathbf{A}})$ where $\theta$ is abelian. Also let $\alpha=(0:\theta)$. Then $D({\mathbf{A}},\theta)$ denotes the algebra ${\mathbf{A}}(\theta)/\Delta_{\theta,\alpha}$.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is an abelian congruence of ${\mathbf{A}}$, and $\alpha=(0:\theta)$. Let $\varphi:=\overline{\alpha}/\Delta_{\theta,\alpha}$ and $D^{o}=\{(a,a)/\Delta_{\theta,\alpha}\,:\,\text{$a\in A$}\}$. Then $\varphi$ is an abelian congruence of $D({\mathbf{A}},\theta)$, $(0:\varphi)=\varphi$, and there exist a surjective homomorphism $h:{\mathbf{A}}(\theta)\rightarrow D({\mathbf{A}},\theta)$ and an isomorphism $h^{\ast}:{\mathbf{A}}/\alpha\cong D({\mathbf{A}},\theta)/\varphi$ such that:
@@
    $D^{o}$ is a transversal for $\varphi$ and $D^{o}=\{0_{C}\,:\,\text{$C\in A/\alpha$}\}\leq D({\mathbf{A}},\theta)$.
@@
    $h$ and $h^{\ast}$ are compatible in the following sense: for all $(a,b)\in\theta$, $h(a,b)/\varphi=h^{\ast}(a/\alpha)$.
@@
That $\varphi$ is abelian and $(0:\varphi)=\varphi$ follows from Theorem 4.6(2). Let $\Delta=\Delta_{\theta,\alpha}$. For each $\alpha$-class $C$ observe that $0_{C}=\{(c,c)\,:\,\text{$c\in C$}\}$ is a $\Delta$-class by Theorem 4.6(1). Hence $D^{o}=\{0_{C}\,:\,\text{$C\in A/\alpha$}\}$. It is easy to check that $D^{o}\leq D({\mathbf{A}},\theta)$. Given $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(a,a)$ and so $(a,b)/\Delta\stackrel{{\scriptstyle\varphi}}{{\equiv}}(a,a)/\Delta\in D^{o}$. Thus $D^{o}$ meets every $\varphi$-class. Suppose $0_{C},0_{C^{\prime}}\in D^{o}$ (so $C,C^{\prime}\in A/\alpha$) and $0_{C}\stackrel{{\scriptstyle\varphi}}{{\equiv}}0_{C^{\prime}}$. Pick $c\in C$ and $c^{\prime}\in C^{\prime}$. Then $(c,c)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(c^{\prime},c^{\prime})$, so $(c,c^{\prime})\in\alpha$, so $C=C^{\prime}$. Thus $D^{o}$ is a transversal for $\varphi$.
Let $h:{\mathbf{A}}(\theta)\rightarrow D({\mathbf{A}},\theta)$ be the natural homomorphism $(a,b)\mapsto(a,b)/\Delta$. To prove item (2), we must show that for all $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(c,c)$ for some $c\in A$ if and only if $a=b$; and this follows from Theorem 4.6(1). Finally, we have obvious isomorphisms ${\mathbf{A}}/\alpha\cong{\mathbf{A}}(\theta)/\overline{\alpha}$ and ${\mathbf{A}}(\theta)/\overline{\alpha}\cong({\mathbf{A}}(\theta)/\Delta)/(\overline{\alpha}/\Delta)=D({\mathbf{A}},\theta)/\varphi$, and we take $h^{\ast}$ to be their composition: $h^{\ast}(a/\alpha)=((a,a)/\Delta)/\varphi$. It is easy to check that item (3) is satisfied.

### v2

@@ 3. Tools
In the context of the proof of Theorem 3.9, it can be further shown that the finite field $\mathbb{F}$ acts on $C_{1}\times\cdots\times C_{k}$ in a coordinate-wise fashion, meaning that each $\theta$-class $C_{i}$ can be viewed as a vector space over $\mathbb{F}$. In congruence modular varieties, Freese proved an analogous result when $\mathbf{A}$} is infinite: there is a division ring $\mathbb{F}$ determined by $\mathbf{A}$} and the abelian minimal congruence $\theta$ such that every $\theta$-class inherits the structure of a left $\mathbb{F}$-vector space. In section 6 we will extend Freese’s result to arbitrary algebras in varieties with a weak difference term, via a new method for constructing the division ring and its action.
@@
Suppose $\mathbf{A}$} is an algebra with a weak difference term and $\theta\in\operatorname{Con}({\mathbf{A}})$ is abelian. Let $D_{1},D_{2}$ be subuniverses of $\mathbf{A}$} which are transversals for $\theta$. Then there exists an automorphism $\sigma$ of $\mathbf{A}$} satisfying $\sigma(D_{1})=D_{2}$ and $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}x$ for all $x\in A$.
@@
Observe first that $\sigma(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}\tau(x)\stackrel{{\scriptstyle\theta}}{{\equiv}}d(x,x,x)=x$ for all $x\in A$. Fix a $\theta$-class $C$; by construction, there exist $b,c\in C$ such that $\pi_{1}(x)=b$ and $\pi_{2}(x)=c$ for all $x\in C$. Pick an abelian group operation ++ on $C$ such that d(x,y,z)=xy+zd(x,y,z)=x-y+z for all $x,y,z\in C$. Then $\sigma(x)=x-b+c$ and $\tau(x)=x-c+b$ for all $x\in C$, and since this is true for every $\theta$-class, we get $\sigma\circ\tau=\tau\circ\sigma=\mathrm{id}_{A}$, proving $\sigma$ is a bijection. Our aim is to prove that $\sigma$ is the desired automorphism.
Let $f$ be an $n$-ary basic operation. Let $a_{1},\ldots,a_{n}\in A$ and put $a=f(a_{1},\ldots,a_{n})$; we need to prove $\sigma(a)=f(\sigma(a_{1}),\ldots,\sigma(a_{n}))$. Let $b_{i}=\pi_{1}(a_{i})$ and $c_{i}=\pi_{2}(a_{i})$, and note that $a_{i},b_{i},c_{i}$ belong to a common $\theta$-class $C_{i}$ for each $i=1,\ldots,n$. Let $b=f(b_{1},\ldots,b_{n})$ and $c=f(c_{1},\ldots,c_{n})$, and note that $b\in D_{1}$ and $c\in D_{2}$ (since $D_{1},D_{2}\leq{\mathbf{A}}$), so $\pi_{1}(a)=b$ and $\pi_{2}(a)=c$. Thus what must be shown is
@@
This is a special case of [16, Theorem 6.16]. ∎
@@
(1) follows from [16, Theorem 3.27] as corrected in [17]. (2) is a special case of [16, Theorem 6.24]. ∎
Our last result in this section generalizes [13, Lemma 2.6].
@@
$\delta\wedge\tau=\sigma$ implies $C(\delta,\tau;\sigma)$ by e.g. [16, Theorem 2.19(8)], proving $(\sigma:\tau)\geq\delta$. Thus to prove $(\delta:\varepsilon)=(\sigma:\tau)$, it is enough to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\sigma)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. We have $C(\alpha,\tau;\sigma)\iff C(\alpha,\tau;\delta)$ by [16, Theorem 2.19(4)], since $\tau\wedge\sigma=\tau\wedge\delta$. Thus we only need to show $C(\alpha,\varepsilon;\delta)\iff C(\alpha,\tau;\delta)$. The forward implication follows immediately by monotonicity. Conversely, suppose $C(\alpha,\tau;\delta)$. Then $C(\alpha,\delta\circ\tau\circ\delta;\delta)$ by [16, Theorem 2.19(3)], i.e., $C(\alpha,\varepsilon;\delta)$.
(2) If $\varepsilon/\delta$ is abelian, then we can apply item (1). Assume $\varepsilon/\delta$ is nonabelian. Then $\tau/\sigma$ is also nonabelian by Proposition 3.13(1). As in the proof of (1), we only need to prove $C(\alpha,\tau;\delta)\implies C(\alpha,\varepsilon;\delta)$ for all $\alpha\in\operatorname{Con}({\mathbf{A}})$ with $\alpha\geq\delta$. Assume $C(\alpha,\tau;\delta)$. If $\alpha\geq\varepsilon$, then $C(\tau,\tau;\delta)$ by monotonicity, which contradicts nonabelianness of $\tau/\sigma$. Thus $\alpha\ngeq\varepsilon$. Since $\alpha\geq\delta$ and $\delta\prec\varepsilon$, we get $\alpha\wedge\varepsilon=\delta$ and hence $C(\alpha,\varepsilon;\delta)$. ∎
4. The difference algebra of an abelian congruence
In this section we revisit an old construction of Hagemann and Herrmann [10] and Freese [5] in congruence modular varieties. We will show that the construction adapts nicely to varieties with a weak difference term. In the following definition, we follow the notation of [6].
@@ 4. The difference algebra of an abelian congruence
    $\mbox{If~}\begin{pmatrix}a&c\\ b&d\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi),~\mbox{then}~\begin{pmatrix}b&d\\ a&c\end{pmatrix}\in\Delta_{\mathsf{h}}(\theta,\varphi).\\$
@@
  4. (4)
    If $\mathbf{A}$} has a weak difference term and $\theta$ is abelian, then $\Delta_{\mathsf{h}}(\theta,\theta)=M(\theta,\theta)$.
@@
(1) $M_{\theta,\varphi}$ is the reflexive, symmetric subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$ generated by $\{((a,a),(b,b))\,:\,\text{$(a,b)\in\varphi$}\}$.
(2) This follows from (1) and the fact that $M(\theta,\varphi)$ is vertically symmetric.
(3) We can consider $\Delta_{\mathsf{h}}(\theta,\varphi)$ as a binary relation on the set of its rows. As such, it is a reflexive subuniverse of ${\mathbf{A}}(\varphi)\times{\mathbf{A}}(\varphi)$, and if $\theta$ is abelian, a weak difference term for $\mathbf{A}$}, interpreted coordinatewise in ${\mathbf{A}}(\varphi)$, will satisfy the hypothesis of Lemma 3.1 for this relation.
(4) Let $\rho=M_{\theta,\theta}$; then $\rho$ is a reflexive subuniverse of ${\mathbf{A}}(\theta)\times{\mathbf{A}}(\theta)$. If $\theta$ is abelian and $\mathbf{A}$} has a weak difference term $d$, then $d$ interpreted coordinatewise in ${\mathbf{A}}(\theta)$ satisfies the hypothesis of Lemma 3.1 for $\rho$. Thus $\rho$ is transitive by Lemma 3.1, implying $\Delta_{\theta,\theta}=M_{\theta,\theta}$ by (1) and hence $\Delta_{\mathsf{h}}(\theta,\theta)=M(\theta,\theta)$. ∎
@@
The next theorem has its roots in [10, Proposition 1.1].
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is a nonzero abelian congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. Let $\Delta=\Delta_{\theta,\alpha}$, let $\eta_{1},\eta_{2},\overline{\theta}$ be as in Definition 4.2, and let $\varepsilon=\overline{\theta}\wedge\Delta$. Then:
@@
$\scriptstyle{\eta_{1}}$$\scriptstyle{\eta_{2}}$$\scriptstyle{0}$$\scriptstyle{1=\overline{1}}$$\scriptstyle{\overline{\theta}}$$\operatorname{Con}({\mathbf{A}}(\theta))=$$\operatorname{Con}({\mathbf{A}})=$$\scriptstyle{0}$$\scriptstyle{\theta}$$\scriptstyle{1}$$\scriptstyle{\alpha}$$\scriptstyle{\overline{\alpha}}$$\scriptstyle{\Delta}$$\scriptstyle{\varepsilon}$
@@
As $\overline{\theta}$ is abelian, we use Proposition 3.11(2) to get $C(\overline{\theta},\overline{\theta};\varepsilon)$, i.e., $\overline{\theta}/\varepsilon$ is abelian. As $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get that $\overline{\alpha}/\Delta$ is abelian by Proposition 3.13(1). Since also $\overline{\theta}/\eta_{1}$ is abelian and $(\eta_{1},\overline{\theta})\searrow(0,\eta_{2})\nearrow(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, we get $(\eta_{1}:\overline{\theta})=(0:\eta_{2})=(\varepsilon:\overline{\theta})=(\Delta:\overline{\alpha})$ by Proposition 3.14(1). Finally, $(\eta_{1}:\overline{\theta})=\overline{(0:\theta)}=\overline{\alpha}$ by Lemma 2.3(2) and the fact that $(0:\theta)=\alpha$, so $(\Delta:\overline{\alpha})=\overline{\alpha}$. This completes the proof of (2).
Now assume $0\prec\theta$. Then $\eta_{i}\prec\overline{\theta}$ for $i=1,2$. Since $\eta_{1}\wedge\eta_{2}=0$, we get $0\prec\eta_{1},\eta_{2}$ (see e.g. [24, Corollary 2.29(ii)]), and every maximal chain in the interval $I[0,\overline{\theta}]$ has length 2 [24, Theorem 2.37]. Thus $I[0,\overline{\theta}]$ has height 2. Note that this also implies $0\prec\varepsilon\prec\overline{\theta}$. Because $\overline{\theta}/\varepsilon$ is abelian and $(\varepsilon,\overline{\theta})\nearrow(\Delta,\overline{\alpha})$, the map $J:I[\varepsilon,\overline{\theta}]\rightarrow I[\Delta,\overline{\alpha}]$ given by $J(\gamma)=\gamma\vee\varepsilon$ is surjective by Proposition 3.13(2). Thus $\Delta\prec\overline{\alpha}$. Now let $\varphi\in\operatorname{Con}({\mathbf{A}}(\mu))$ with $\Delta\leq\varphi$ but $\overline{\alpha}\nleq\varphi$. Then $\overline{\alpha}\wedge\varphi=\Delta$, so $C(\varphi,\overline{\alpha};\Delta)$, so $\varphi\leq(\Delta:\overline{\alpha})=\overline{\alpha}$, so $\varphi=\overline{\alpha}\wedge\varphi=\Delta$. This proves (6). ∎
The next definition is essentially from [5, 6] in the congruence modular setting; we extend it to varieties with a weak difference term.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, and $\theta\in\operatorname{Con}({\mathbf{A}})$ where $\theta$ is abelian. Let $\alpha=(0:\theta)$.
  1. (1)
    The difference algebra for $\theta$, denoted $D({\mathbf{A}},\theta)$, is the algebra ${\mathbf{A}}(\theta)/\Delta_{\theta,\alpha}$.
  2. (2)
    The derived congruence, which we denote by $\varphi$, is the congruence $\overline{\alpha}/\Delta_{\theta,\alpha}$.
  3. (3)
    The canonical transversal is the set $D^{o}:=\{0_{E}\,:\,\text{$E\in A/\alpha$}\}$.
@@
Suppose $\mathcal{V}$ is a variety with a weak difference term, ${\mathbf{A}}\in\mathcal{V}$, $\theta$ is an abelian congruence of $\mathbf{A}$}, and $\alpha=(0:\theta)$. Let $D({\mathbf{A}},\theta)$ be the difference algebra for $\theta$, let $\varphi$ be the derived congruence, let $D^{o}$ be the canonical transversal, and let $\nu:{\mathbf{A}}(\theta)\rightarrow D({\mathbf{A}},\theta)$ be the natural projection map. Then:
@@
    $\varphi$ is an abelian congruence of $D({\mathbf{A}},\theta)$.
@@
    $D^{o}$ is a transversal for $\varphi$ and $D^{o}\leq D({\mathbf{A}},\theta)$.
  4. (4)
    $\nu^{-1}(D^{o})=0_{A}$.
  5. (5)
    There exists an isomorphism $h:{\mathbf{A}}/\alpha\cong D({\mathbf{A}},\theta)/\varphi$ such that for all $(a,b)\in\theta$, $\nu(a,b)/\varphi=h(a/\alpha)$.
@@
Items (1) and (2) follow from Theorem 4.6(2). Let $\Delta=\Delta_{\theta,\alpha}$. For each $\alpha$-class $E$ observe that $0_{E}=\{(c,c)\,:\,\text{$c\in E$}\}$ is a $\Delta$-class by Theorem 4.6(1). Hence $D^{o}\subseteq D({\mathbf{A}},\theta)$. It is easy to check that $D^{o}\leq D({\mathbf{A}},\theta)$. Given $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(a,a)$ and so $(a,b)/\Delta\stackrel{{\scriptstyle\varphi}}{{\equiv}}(a,a)/\Delta\in D^{o}$. Thus $D^{o}$ meets every $\varphi$-class. Suppose $0_{E},0_{E^{\prime}}\in D^{o}$ and $0_{E}\stackrel{{\scriptstyle\varphi}}{{\equiv}}0_{E^{\prime}}$. Pick $c\in E$ and $c^{\prime}\in E^{\prime}$. Then $(c,c)\stackrel{{\scriptstyle\overline{\alpha}}}{{\equiv}}(c^{\prime},c^{\prime})$, so $(c,c^{\prime})\in\alpha$, so $E=E^{\prime}$. Thus $D^{o}$ is a transversal for $\varphi$.
To prove item (4), we must show that for all $(a,b)\in\theta$ we have $(a,b)\stackrel{{\scriptstyle\Delta}}{{\equiv}}(c,c)$ for some $c\in A$ if and only if $a=b$; and this follows from Theorem 4.6(1). To prove item (5), define $h$ by $h(a/\alpha)=((a,a)/\Delta)/\varphi$.